# UniCtrl: Improving the Spatiotemporal Consistency of Text-to-Video Diffusion Models via Training-Free Unified Attention Control

### v3

@@ 4.4 Cross-Frame Unified Attention Control
As illustrated in Figure 3, we integrate Cross-Frame Self-Attention Control, Motion Injection and Spatiotemporal Synchronization into a cohesive framework termed Cross-Frame Unified Attention Control. In the output branch, we employ the key $K^{0}$ and value $V^{0}$ derived from the initial frame to maintain cross-frame semantic consistency. A separate branch is utilized to preserve the query specifically for motion control, replacing the output branch’s query with $Q_{\text{m}}$ from the motion control branch. To prevent spatiotemporal inconsistencies, we synchronize the latent representation of the motion branch with the output branch’s preceding latent state before each sampling step. This approach is further detailed and formalized presented in Algorithm 2.
@@ 5.1 Experimental Setup
To evaluate the effectiveness of our model, we collect prompts from two datasets UCF-101[53] and MSR-VTT[65] for generating videos. Following Ge et al[8], we use the same UCF-101 prompts for our experiments. We also randomly select 100 unique prompts from the MSR-VTT dataset for our evaluation. Those two parts of data consist of our dataset for evaluations. Next we briefly introduce evaluation metrics and we will provide details in the supplementary material.
5.1.1 Metric
To quantitatively evaluate our results, we consider standard metrics following [46, 63]:
$\bullet$ DINO: To evaluate the spatiotemporal consistency in the generated video, we employ DINO[38] to compare the cosine similarity between initial frame and subsequent frames. In our experiments, we utilize the DINO-vits16[5] model to compute the DINO cosine similarity.
$\bullet$ RAFT: To compare the magnitude of motion in the videos, we utilize RAFT[54] to estimate the optical flow, thereby inferring the degree of motion. We utilize the off-the-shelf RAFT model from torchvision [34].
5.1.2 Backbones
@@
$\bullet$ AnimateDiff [16] introduces a practical framework for adding motion dynamics to personalized text-to-image models, such as those created by Stable Diffusion, without the need for model-specific adjustments. Central to AnimateDiff is a motion module, trainable once and universally applicable across personalized text-to-image models derived from the same base model, leveraging transferable motion priors from real-world videos for animation.
$\bullet$ VideoCrafter [6] introduces two novel diffusion models for video generation: T2V, which synthesizes high-quality videos from text inputs, achieving cinematic
@@
-quality resolutions up to 1024$\times$576, and I2V, the first open-source model that transforms images into videos while preserving content, structure, and style. Both models represent significant advancements in open-source video generation technology, offering new tools for researchers and engineers. We use T2V version in our experiments.
5.1.3 Baseline
@@
We select FreeInit [63] as our baseline since FreeInit is a training free method that attempts to improve subject appearance and temporal consistency of generated videos through iteratively refining the spatial-temporal low-frequency components of the initial latent during inference. However, given that our method UniCtrl operates on the attention mechanism and FreeInit adjusts the frequency domain of the latent space, UniCtrl and FreeInit are orthogonal approaches. Both are training-free methods capable of enhancing the spatiotemporal consistency of generated videos via diffusion models. We demonstrate the integration of UniCtrl and FreeInit both in 5.2 and 5.3.
5.2 Quatitative Comparisons
@@
Figure 4: Qualitative Comparisons. We showcase UniCtrl’s adaptability to varied text prompts, leveraging UniCtrl to significantly improve temporal consistency and preserve motion diversity. Comparative inference results with FreeInit are presented for context. Additionally, we demonstrate UniCtrl’s seamless integration with FreeInit.
@@
Qualitative comparisons, depicted in Figure 4, reveal that our UniCtrl method markedly improves spatiotemporal consistency and maintains motion diversity. For example, with the text prompt "walking with a dog", FreeInit produces inconsistent appearances for both the lady and the dog, whereas UniCtrl ensures consistent representations of both entities. Furthermore, when processing the prompt “A young woman walks through flashing lights", UniCtrl maintains the detailed features of the young woman’s dress while ensuring her walking motion remains natural, in contrast to the vanilla AnimateDiff model. Additionally, we demonstrate UniCtrl’s flawless integration with FreeInit, enhancing motion in the background while consistently preserving the young woman’s appearance. Additional qualitative findings can be found in the supplementary material.
5.3 Quanlitative Comparisons
@@
Table 1: Quantitative Comparisons on UCF-101 and MSR-VTT. UniCtrl significantly improves the temporal consistency while keeps the motion in the generated videos. $c$ indicates the motion injection degree. $I$ indicates number of iteration for FreeInit.
@@
AnimateDiff [16] 93.99 31.38
FreeInit + AnimateDiff ($I=3$) 96.15 14.79
UniCtrl + AnimateDiff ($c=1$) 96.34${}_{(+00.19)}$ 25.70${}_{(+10.91)}$
VideoCrafter [6] 93.39 28.25
FreeInit + VideoCrafter ($I=3$) 96.61 6.90
UniCtrl + VideoCrafter ($c=1$) 94.13${}_{(-02.48)}$ 28.78${}_{(+21.88)}$
UniCtrl + FreeInit + AnimateDiff ($I=3$, $c=1$) 96.48 17.91
For quantitative comparisons, the quantitative results on UCF-101 and MSR-VTT are reported in Table 1. We compare the backbones and backbones augmented by UniCtrl and FreeInit respectively. According to the metrics, UniCtrl significantly improves the spatiotemporal consistency in the generated videos across all backbones on both prompt sets from $0.74$ to $2.35$. While FreeInit achieves remarkable improvements over spatiotemporal consistency, we found that UniCtrl outperforms FreeInit on the strength of motions on both AnimateDiff and VideoCrafter by a large margin from $10.91$ to $21.88$. Note that we purposely choose the motion injection degree $c=1$ to demonstrate how UniCtrl can preserve the motion compare with FreeInit. However, there still exists a trade off between spatiotemporal consistency and motion diversity. In section 5.4, we show spatiotemporal consistency can be further improved with a smaller motion injection degree. Thus, we recommend motion injection degree $c=0.4$ for real-world applications. Lastly, we showcase how UniCtrl and FreeInit can improve AnimateDiff together and we found this integration can improve both spatiotemporal consistency and motion diversity. We will introduce the details how we integrate UniCtrl and FreeInit in the supplementary material. Note that we obtained different score for FreeInit because we randomly sampled a different set of prompts from MSR-VTT and we used our own implementations for FreeInit on VideoCrafter since we cannot find official implementation.
@@ 5.4 Ablation Study
Method DINO ($\uparrow$) RAFT
AnimateDiff [16] 93.99 31.81
UniCtrl w/o SAC + AnimateDiff 93.99 31.81
UniCtrl w/o MI + AnimateDiff 97.98 4.19
UniCtrl w/o SS + AnimateDiff 93.98 31.78
UniCtrl ($c=0)$ + AnimateDiff 97.98 4.19
UniCtrl ($c=0.2)$ + AnimateDiff 97.36 9.73
UniCtrl ($c=0.4)$ + AnimateDiff 96.62 16.15
UniCtrl ($c=0.6)$ + AnimateDiff 96.41 22.14
UniCtrl ($c=0.8)$ + AnimateDiff 96.34 24.60
UniCtrl ($c=1.0)$ + AnimateDiff 96.34 25.70
We assess the impact of each UniCtrl module, the efficacy of motion injection degree, and corroborate our design choice of swapping Key and Value together in the SAC procedure through the subsequent ablation studies.
5.4.1 Evaluating the Impact of SAC, MI, and SS
To assess the contribution of the SAC, we conducted experiments on both datasets using AnimateDiff as the baseline, incorporating our pipeline but disabling SAC. For motion injection, we set motion injection degree to be $1$. The findings reveal our method without SAC works exactly the same the backbone because the motion branch is the exactly the same as the output branch. This observation underscores the critical role of SAC in enhancing spatiotemporal consistency, as evidenced by the comparisons with the scores fom the vanilla AnimateDiff.
@@ 5.4.1 Evaluating the Impact of SAC, MI, and SS
Finally, the necessity of the Stabilization Strategy (SS) was examined by ex
cluding SS from our pipeline and conducting experiments across both datasets, again using AnimateDiff as the reference point and setting motion injection degree to be $1$. The outcomes showed diminished spatiotemporal consistency in comparison to the baseline, while motion diversity was not significantly impacted. These results emphasize the importance of integrating SS into our pipeline, as corroborated by the UniCtrl’s findings presented in Table 1, illustrating the essential role of SS in achieving the desired balance of spatiotemporal consistency and motion diversity.
5.4.2 Influence of Motion Injection Degree
@@
5.4.3 Influence of Swapping Key and Value
Figure 5: Each row displays a sequence of video frames generated from the identical prompt: A panda standing on a surfboard in the ocean under moonlight. The section labeled K and V mismatch illustrates the frames produced when there is a discrepancy between key and value pairs. Conversely, the section titled K and V match showcases frames generated when key and value pairs are in alignment.
We aim to provide a qualitative example to underscore the importance of simultaneously modifying both key and value, as highlighted in our discussion on the impact of key and value mismatches in Section 4.1. Initially, we alter only the value while maintaining the same key within the Spatial Attention Component (SAC) performing the UniCtrl pipeline. As depicted in the first row of Fig. 5, the panda begins to fade in the subsequent frames, indicating a substantial decrease in the quality of the generated videos with respect to spatiotemporal consistency. This decline is particularly evident when compared to the approach of modifying both the key and value concurrently using the same UniCtrl pipeline.

### v4

@@ 4.4 Cross-Frame Unified Attention Control
As illustrated in Figure 3, we integrate cross-frame self-attention control, motion injection and spatiotemporal synchronization into a cohesive framework termed Cross-Frame Unified Attention Control(UniCtrl). In the output branch, we employ the key $K^{0}$ and value $V^{0}$ derived from the initial frame to maintain cross-frame semantic consistency. A separate branch is utilized to preserve the query specifically for motion control, replacing the output branch’s query with $Q_{\text{m}}$ from the motion control branch. To prevent spatiotemporal inconsistencies, we synchronize the latent representation of the motion branch with the output branch’s preceding latent state before each sampling step.
@@ 5.1 Experimental Setup
To evaluate the effectiveness of our model, we collect prompts from two datasets UCF-101 (Soomro et al., 2012) and MSR-VTT (Xu et al., 2016) for generating videos. Following Ge et al(Ge et al., 2023a), we use the same UCF-101 prompts for our experiments. We also randomly selected 100 unique prompts from the MSR-VTT dataset for our evaluation. Those two parts of data consist of our dataset for evaluation. To mitigate the stochasticity inherent in diffusion models, all experiments were conducted using multiple random seeds. The scores reported in the tables represent the average results across these runs, and we provide the corresponding standard deviations in the Appendix. Next, we briefly introduce evaluation metrics and we also provide details in the Appendex.
Metric
To quantitatively evaluate our results, we consider standard metrics following (Singer et al., 2022; Wu et al., 2023):
$\bullet$ DINO: To evaluate the spatiotemporal consistency in the generated video, we employ DINO(Oquab et al., 2024) to compare the cosine similarity between the initial frame and subsequent frames. In our experiments, we utilize the DINO-vits16(Caron et al., 2021) model to compute the DINO cosine similarity.
$\bullet$ RAFT: To compare the magnitude of motion in the videos, we utilize RAFT(Teed & Deng, 2020) to estimate the optical flow, thereby inferring the degree of motion. We utilize the off-the-shelf RAFT model from torchvision (maintainers & contributors, 2016).
$\bullet$ FVD: To assess video quality, particularly the quality of individual frames during realistic motions, we employ the Content-Debiased FVD metric (Ge et al., 2024) to evaluate the generated videos. In our experiments, we use the official library to compute the FVD.
$\bullet$ FVMD: To evaluate the quality of motion consistency in video generation, we use the Fréchet Video Motion Distance (FVMD) metric (Liu et al., 2024) to assess temporal coherence based on velocity and acceleration patterns. In our experiments, we employ the official library to calculate the FVMD with the ground truth videos from our prompt dataset.
Backbones
@@
$\bullet$ AnimateDiff (Guo et al., 2023b) introduces a practical framework for adding motion dynamics to personalized text-to-image models, such as those created by Stable Diffusion, without the need for model-specific adjustments. Central to AnimateDiff is a motion module, trainable once and universally applicable across personalized text-to-image models derived from the same base model, leveraging transferable motion priors from real-world videos for animation.
@@
$\bullet$ VideoCrafter (Chen et al., 2023) introduces two novel diffusion models for video generation: T2V, which synthesizes high-quality videos from text inputs, achieving cinematic-quality resolutions up to 1024$\times$576, and I2V, the first open-source model that transforms images into videos while preserving content, structure, and style. Both models represent significant advancements in open-source video generation technology, offering new tools for researchers and engineers. We use the T2V version in our experiments.
$\bullet$ AnimateLCM (Wang et al., 2024) introduces a novel framework for efficient and high-quality video generation by leveraging consistency learning. It builds upon the principles of the Consistency Model (CM) and Latent Consistency Model (LCM) from image diffusion models, accelerating video generation with minimal steps. AnimateLCM employs a decoupled consistency learning strategy, separating the learning of image generation priors and motion generation priors.
Baseline
@@
We select FreeInit (Wu et al., 2023) as our baseline since FreeInit is a training-free method that attempts to improve the subject appearance and temporal consistency of generated videos through iteratively refining the spatial-temporal low-frequency components of the initial latent during inference. However, given that our method UniCtrl operates on the attention mechanism and FreeInit adjusts the frequency domain of the latent space, UniCtrl and FreeInit are orthogonal approaches. Both are training-free methods capable of enhancing the spatiotemporal consistency of generated videos via diffusion models. We demonstrate the integration of UniCtrl and FreeInit both in 5.2 and 5.3.
5.2 Qualitative Comparisons
@@
Figure 4: Qualitative Comparisons. We demonstrate UniCtrl’s adaptability to diverse prompts, enhancing temporal consistency and preserving motion diversity. Comparative inference results with FreeInit are presented for context. Additionally, we demonstrate UniCtrl’s seamless integration with FreeInit.
@@
Qualitative comparisons, depicted in Figure 4, reveal that our UniCtrl method markedly improves spatiotemporal consistency and maintains motion diversity. For example, with the text prompt “walking with a dog”, FreeInit produces inconsistent appearances for both the lady and the dog, whereas UniCtrl ensures consistent representations of both entities. Furthermore, when processing the prompt “A young woman walks through flashing lights”, UniCtrl maintains the detailed features of the young woman’s dress while ensuring her walking motion remains natural, in contrast to the vanilla AnimateDiff model. Additionally, we demonstrate UniCtrl’s flawless integration with FreeInit, consistently preserving the young woman’s appearance.
We show additional qualitative results in Figure 5 and Figure 6 to demonstrate the efficacy of UniCtrl to significantly improving spatiotemporal consistency and preserve motion dynamics for a diverse set of prompts.
Figure 5: We provide additional qualitative examples across various backbones to demonstrate UniCtrl’s capability in enhancing spatiotemporal consistency while effectively preserving motion dynamics.
Figure 6: We present more qualitative examples to show UniCtrl’s ability to improve spatiotemporal consistency and significantly preserve motion dynamics.
5.3 Quantitative Comparisons
@@
Table 1: Quantitative Comparisons on UCF-101 and MSR-VTT. UniCtrl significantly improves the temporal consistency while keeping the motion in the generated videos. $c$ indicates the motion injection degree. $I$ indicates the number of iterations for FreeInit.
@@
AnimateDiff (Guo et al., 2023b) 94.26 25.44
FreeInit + AnimateDiff ($I=3$) 96.04 11.62
UniCtrl + AnimateDiff ($c=1$) 96.38 ($+$00.34) 23.29 ($+$11.67)
VideoCrafter (Chen et al., 2023) 93.53 29.20
FreeInit + VideoCrafter ($I=3$) 96.75 7.46
UniCtrl + VideoCrafter ($c=1$) 95.55 (--01.20) 25.11 ($+$17.65)
AnimateLCM (Wang et al., 2024) 92.96 20.80
FreeInit + AnimateLCM ($I=3$) 94.90 14.64
UniCtrl + AnimateLCM ($c=1$) 95.40 ($+$00.50) 17.53 ($+$2.89)
UniCtrl + FreeInit + AnimateDiff ($I=3$, $c=1$) 96.85 9.82
For quantitative comparisons, the quantitative results on UCF-101 and MSR-VTT are reported in Table 1. We compare the backbones and backbones augmented by UniCtrl and FreeInit respectively. According to the metrics, UniCtrl significantly improves the spatiotemporal consistency in the generated videos across all backbones on both prompt sets from $2.12$ to $2.44$. While FreeInit achieves remarkable improvements over spatiotemporal consistency, we found that UniCtrl outperforms FreeInit on the strength of motions on both AnimateDiff and VideoCrafter by a large margin from $11.67$ to $17.65$. Note that we purposely chose the motion injection degree $c=1$ to demonstrate how UniCtrl can preserve the motion compared with FreeInit. However, there still exists a trade-off between spatiotemporal consistency and motion diversity. In section 5.4, we show spatiotemporal consistency can be further improved with a smaller motion injection degree. Thus, we recommend motion injection degree $c=0.4$ for real-world applications. Lastly, we showcase how UniCtrl and FreeInit can improve AnimateDiff together and we found this integration can improve spatiotemporal consistency together. We will introduce the details of how we integrate UniCtrl and FreeInit in the Appendix. Note that we obtained different scores for FreeInit because we randomly sampled a different set of prompts from MSR-VTT and we used our own implementations for FreeInit on VideoCrafter since we cannot find official implementation.
Table 2: Quality results on UCF-101, which show a significant improvement in both video quality and motion consistency in video generation by UniCtrl.
Method FVD ($\downarrow$) FVMD ($\downarrow$)
AnimateDiff (Guo et al., 2023b) 1069.90 27124.17
FreeInit + AnimateDiff 958.97 25078.05
UniCtrl + AnimateDiff 819.74 8864.07
Furthermore, regarding video quality and motion consistency, we present the quality results on UCF-101 in Table 2. We compare AnimateDiff with FreeInit augmented and Unictrl augmented. The metrics show that UniCtrl significantly enhances the quality of individual video frames, improving the score from 1069.90 to 819.74. Additionally, we observe a substantial improvement in FVMD, decreasing from 27,124.17 to 8,864.07, which indicates a marked enhancement in motion consistency. These results further demonstrate our model’s capability to produce higher-quality videos with improved motion coherence, highlighting the significant effectiveness of our approach in advancing text-to-video models.
@@ 5.4 Ablation Study
In this section, we assess the impact of each UniCtrl module and the efficacy of the motion injection degree. We further corroborate our design choice of swapping Key and Value together in the SAC procedure through the subsequent ablation studies.
The Impact of SAC, MI, and SS
@@
Method DINO ($\uparrow$) RAFT ($\uparrow$)
AnimateDiff (Guo et al., 2023b) 94.26 25.44
UniCtrl w/o SAC + AnimateDiff 94.26 25.42
UniCtrl w/o MI + AnimateDiff 98.08 4.12
UniCtrl w/o SS + AnimateDiff 94.26 21.90
only SAC + AnimateDiff 98.08 4.12
only MI + AnimateDiff 94.26 25.42
only SS + AnimateDiff 94.26 25.42
UniCtrl ($c=0)$ + AnimateDiff 98.08 4.12
UniCtrl ($c=0.2)$ + AnimateDiff 97.41 9.04
UniCtrl ($c=0.4)$ + AnimateDiff 96.69 15.56
UniCtrl ($c=0.6)$ + AnimateDiff 96.46 20.16
UniCtrl ($c=0.8)$ + AnimateDiff 96.37 22.50
UniCtrl ($c=1.0)$ + AnimateDiff 96.38 23.29
To assess the contribution of the SAC, we conducted experiments on both datasets using AnimateDiff as the baseline, incorporating our pipeline but disabling SAC. For motion injection, we set the motion injection degree to be $1$. The findings reveal our method without SAC works the same on the backbone because the motion branch is exactly the same as the output branch. This observation underscores the critical role of SAC in enhancing spatiotemporal consistency, as evidenced by the comparisons with the scores from the vanilla AnimateDiff.
@@
Finally, the necessity of the Spatiotemporal Synchronization (SS) was examined by excluding SS from our pipeline and conducting experiments across both datasets, again using AnimateDiff as the reference point and setting motion injection degree to be $1$. The outcomes showed diminished spatiotemporal consistency in comparison to the baseline, while motion diversity was not significantly impacted. These results emphasize the importance of integrating SS into our pipeline, as corroborated by the UniCtrl’s findings presented in Table 1, illustrating the essential role of SS in achieving the desired balance of spatiotemporal consistency and motion diversity.
Additionally, we conduct experiments on each individual component to provide a deeper understanding of their respective roles and interactions, as shown in Table 3. The results align with our previous analysis: when only SAC is present, we observe a significant increase in the DINO score, confirming the critical role of SAC in enhancing spatiotemporal consistency. In contrast, when only motion injection is applied without SAC and the consistency SAC brings, motion injection alone proves ineffective. Similarly, SS alone shows no impact; only when combined with SAC and MI does it effectively achieve spatiotemporal consistency.
The Impact of Motion Injection Degree
@@
The Impact of Swapping Key and Value
Figure 7: Each row displays a sequence of video frames generated from the identical prompt: “A panda standing on a surfboard in the ocean under moonlight.” The section labeled K and V mismatch illustrates the frames produced when there is a discrepancy between key and value pairs. Conversely, the section titled K and V match showcases frames generated when key and value pairs are in alignment.
We aim to provide a qualitative example to underscore the importance of simultaneously modifying both key and value, as highlighted in our discussion on the impact of key and value mismatches in Section 4.1. Initially, we alter only the value while maintaining the same key within the cross-frame Self-Attention Control (SAC) performing the UniCtrl pipeline. As depicted in the first row of Figure 7, the panda begins to fade in the subsequent frames, indicating a substantial decrease in the quality of the generated videos with respect to spatiotemporal consistency. This decline is particularly evident when compared to the approach of modifying both the key and value concurrently using the same UniCtrl pipeline.
@@ 6.2 Ethics Statement
Acknowledgment
We thank Raven, Tom, Martin, Yichi Zhang, Shengyi Qian for their helpful feedback and advice.
Appendix
Here we present some of the sections left for discussion in the main paper. This supplementary report includes discussions on Details of Metrics, Implementation Details, and more Qualitative Results on the effectiveness of the motion injection degree, ablations of each module in UniCtrl, and how UniCtrl and FreeInit (Wu et al., 2023) can work together. These results further demonstrate UniCtrl’s efficacy in improving spatiotemporal consistency and preserving motion dynamics.
Appendix A Details of Metrics
We consider standard metrics: DINO and RAFT following (Singer et al., 2022; Wu et al., 2023) and we present details of our evaluation metrics:
$\bullet$ DINO: To evaluate the spatiotemporal consistency in the generated video, we employ DINO (Oquab et al., 2024) to compare the cosine similarity between the initial frame and subsequent frames. The average DINO score across all consecutive frames is then used as the video’s overall score. In our experiments, we utilize the DINO-vits16(Caron et al., 2021) model to compute the DINO cosine similarity. We present the details of our implementation in Algorithm 3.
$\bullet$ RAFT: To compare the magnitude of motion in the videos, we utilize RAFT (Teed & Deng, 2020) to estimate the optical flow, thereby inferring the degree of motion. To quantify the motion intensity between every two consecutive frames, we employ the RAFT model to estimate the optical flow for each pixel. We calculate the magnitude of all pixel optical flow using the $l^{2}$ norm, and finally, we obtain the average of all magnitude as the score for every consecutive frame in the video. In this process, we utilize the RAFT model from torchvision (maintainers & contributors, 2016). We present the details of the RAFT implementation in Algorithm 4.
Algorithm 3 DINO Score
1:Input: Generated Video $v$
[…]
4:for $n=1$ to $N$ do
[…]
6:end for
7:$DINO\_SCORE=\textsc{average}(\textrm{scores})$
8:Output: $DINO\_SCORE$
Algorithm 4 RAFT Score
1:Input: Generated Video $v$
[…]
3:for $n=1$ to $N$ do
[…]
6:end for
7:$RAFT\_SCORE=\textsc{average}(\textrm{scores})$ Average score from pairs of frames
8:Output: $RAFT\_SCORE$
Appendix B Implementation Details
B.1 Backbones
Two open-sourced text-to-video models are used as the base models for FreeInit evaluation. For VideoCrafter(Chen et al., 2023), we adopt the VideoCrafter1-base-1024-T2V-model. For AnimateDiff (Guo et al., 2023b), we use the mm-sd-v14 motion module with the Realistic-Vision-V5.1 LoRA model for evaluation. For AnimateLCM Wang et al. (2024), we use the AnimateLCM-sd15-t2v motion module and also Realistic-Vision-V5.1 LoRA model for evaluation.
B.2 Inference Details
Experiments on VideoCrafter is conducted on 512 × 320 spatial scale and 16 frames, while experiments on AnimateDiff is conducted on a video size of 512 × 512, 16 frames. AnimateLCM is also conducted on 512 × 512, 16 frames. During the inference process, we use classifier-free guidance for all experiments including the comparisons and ablation studies, with a constant guidance weight 7.5. All experiments are conducted on Nvidia A10G GPU. We conduct experiments across different GPUs such as Nvidia A10G, Nvidia A40, and Nvidia A100 and we observe different inference results using the same seed while the CUDA (NVIDIA et al., 2020) random algorithm is deterministic. We resolve this behavior by initializing latent using NumPy (Harris et al., 2020). We hope this change will further enhance the reproducibility of our code across different CUDA architectures.
Appendix C Detailed Quantitative Results Tables
Since our experiments are conducted with multiple seeds, we report the mean values and standard deviations in the following table as a supplement to the previous tables
Table 4: Detailed Table for Table 1
Method DINO ($\uparrow$) RAFT ($\uparrow$)
AnimateDiff (Guo et al., 2023b) 94.26 $\pm$ 0.63 25.44 $\pm$ 5.81
FreeInit + AnimateDiff ($I=3$) 96.04 $\pm$ 0.22 11.62 $\pm$ 1.72
UniCtrl + AnimateDiff ($c=1$) 96.38 $\pm$ 0.21 23.29 $\pm$ 2.09
VideoCrafter (Chen et al., 2023) 93.53 $\pm$ 0.27 29.20 $\pm$ 1.16
FreeInit + VideoCrafter ($I=3$) 96.75 $\pm$ 0.07 7.46 $\pm$ 0.59
UniCtrl + VideoCrafter ($c=1$) 95.55 $\pm$ 0.17 25.11 $\pm$ 1.98
AnimateLCM (Wang et al., 2024) 92.96 $\pm$ 0.18 20.80 $\pm$ 5.15
FreeInit + AnimateLCM ($I=3$) 94.90 $\pm$ 0.89 14.64 $\pm$ 4.59
UniCtrl + AnimateLCM ($c=1$) 95.40 $\pm$ 0.27 17.53 $\pm$ 3.47
UniCtrl + FreeInit + AnimateDiff ($I=3$, $c=1$) 96.85 $\pm$ 0.19 9.82 $\pm$ 1.13
Table 5: Detailed Table for Table 2
Method FVD ($\downarrow$) FVMD ($\downarrow$)
AnimateDiff (Guo et al., 2023b) 1069.90 $\pm$ 57.89 27124.17 $\pm$ 3234.76
FreeInit + AnimateDiff 958.97 $\pm$ 40.82 25078.05 $\pm$ 5490.52
UniCtrl + AnimateDiff 819.74 $\pm$ 29.42 8864.07 $\pm$ 577.43
Table 6: Detailed Table for Table 3
Method DINO ($\uparrow$) RAFT ($\uparrow$)
AnimateDiff (Guo et al., 2023b) 94.26 $\pm$ 0.63 25.44 $\pm$ 5.81
UniCtrl w/o SAC + AnimateDiff 94.26 $\pm$ 0.62 25.42 $\pm$ 5.86
UniCtrl w/o MI + AnimateDiff 98.08 $\pm$ 0.47 4.12 $\pm$ 2.33
UniCtrl w/o SS + AnimateDiff 94.26 $\pm$ 1.16 21.90 $\pm$ 2.08
only SAC + AnimateDiff 98.08 $\pm$ 0.47 4.12 $\pm$ 2.33
only MI + AnimateDiff 94.26 $\pm$ 0.62 25.42 $\pm$ 5.86
only SS + AnimateDiff 94.26 $\pm$ 0.62 25.42 $\pm$ 5.86
UniCtrl ($c=0)$ + AnimateDiff 98.08 $\pm$ 0.47 4.12 $\pm$ 2.33
UniCtrl ($c=0.2)$ + AnimateDiff 97.41 $\pm$ 0.10 9.04 $\pm$ 1.37
UniCtrl ($c=0.4)$ + AnimateDiff 96.69 $\pm$ 0.14 15.56 $\pm$ 1.44
UniCtrl ($c=0.6)$ + AnimateDiff 96.46 $\pm$ 0.17 20.16 $\pm$ 1.73
UniCtrl ($c=0.8)$ + AnimateDiff 96.37 $\pm$ 0.19 22.50 $\pm$ 1.90
UniCtrl ($c=1.0)$ + AnimateDiff 96.38 $\pm$ 0.21 23.29 $\pm$ 2.09
Appendix D More Qualitative Results
D.1 Motion Injection Degree
In the main paper, we quantitatively demonstrate that higher degrees of motion injection results in videos with more pronounced motion. Here, we provide a qualitative example in Fig. 8. The motion of the Corgi is significantly more pronounced when comparing videos generated with a higher motion injection degree to those with a lower degree, while the appearance of the Corgi remains consistent across frames in each video.
Figure 8: We present 4 sequences of images inferenced from the same prompt but with different Motion Injection Degree. With larger motion injection degree, corgi in each generated video presents more motion while preserve its appearance.
D.2 Qualitative Ablation Results
Figure 9: The first row demonstrates results generated with UniCtrl, which the vehicle and the road are both consistent across the frames. The second row shows frames generated with SAC and MI. The third row explains frames augmented with SS and MI. The fourth row contains frames shows results with SS and SAC. These comparisons serve as qualitative examples for ablation for each module in UniCtrl.
In the main paper, we conduct quantitative ablation studies of each module and we would like to emphasize the indispensable nature of every component within UniCtrl again by presenting Fig. 9. This figure qualitatively illustrates the ablations through a single set of results. By comparing the first and second rows, it becomes evident that the results in the second-row display inconsistencies across frames, both in terms of the car and the road. Intuitively, the motion branch and output branch keep deviating while denoising. SS effectively bridges the difference between the motion branch and the output branch and significantly improves spatiotemporal consistency as a result. By comparing the first and third rows, we observe the third row’s result shows much more inconsistencies which clearly explains the necessity of incorporating SAC in UniCtrl. Lastly, we show the fourth row’s result contains little motion compared with the first row’s result. Thus, we prove that MI is also one of the indispensable modules of UniCtrl.
D.3 Unification between UniCtrl and FreeInit
We state that UniCtrl and FreeInit explore orthogonal directions to improve the spatiotemporal consistency of video diffusion models. Now we want to first explain how to unify UniCtrl and FreeInit. In the framework of FreeInit, it requires $N$ FreeInit iterations, $N=3$ in our implementation, and we apply UniCtrl during the first iteration. We find the unification between UniCtrl and FreeInit improves the quality of generated videos and we showcase additional qualitative examples in Fig. 10.
Figure 10: We present more qualitative examples in order to show the unification between UniCtrl and FreeInit can improve spatiotemporary consistency and preserve motion dynamics, which further demonstrate that UniCtrl and FreeInit are orthogonal works that both can improve spatiotemporary consistency.